OpenStudy (anonymous):

Solve the linear system using elimination. 8x - (1/2)y = -38 (1/4)x - (1/2)y = -7

OpenStudy (ajprincess):

Subtract the second equation from the first equation and pls tell me vat u get?

OpenStudy (anonymous):

Help me subtract, I'm new to this unit. :(

OpenStudy (ajprincess):

8x-(1/2)y=-38-(1) (1/4)x-(1/2)y=-7-(2) (2)-(1)=>8x-(1/2)y-(1/4)x-(-(1/2)y)=-38-(-7) 8x-(1/2)y-(1/4)x+(1/2)y=-38+7 Does that help @Rubie

OpenStudy (anonymous):

8x - (1/2)y +38=(1/4)x - (1/2)y +7 (31/4)x+31=0 (31/4)x=-31 31x=-31(4) x=-4 then substitute the x with the value -4 and you get the value of y

OpenStudy (anonymous):

? = -31 ? I don't quite get it.

OpenStudy (anonymous):